OpenStudy (anonymous):

two similar pyramids have lateral areas of 20 ft^2 and 45 ft^2. the volume of the smaller pyramid is 8ft^3. what is the volume of the larger pyram

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

areas come with the square, volume with the cube. original ratio is \[\frac{20}{45}=\frac{4}{9}\] the square root of that is \[\frac{2}{3}\] and the cube is \[\frac{8}{27}\] so that is the the ratio of the volume. so take \[8\times \frac{27}{8}=27\] is the answer

OpenStudy (anonymous):

hey hero